# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v3

@@ 3.1. Proofs
$s_{N}\left({\bf p}|{\bf q}\right):=\sum_{i\in\mathcal{I}_{N}}p_{i}\log\frac{q_{i}}{p_{i}},\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
If $p_{i}=0$, then, $p_{i}\log\frac{q_{i}}{p_{i}}:=0$. $s_{N}$ is continuous on $\overline{\mathcal{P}_{N}}$. Then, by Jensen’s inequality, $s_{N}({\bf p}|{\bf q})\leq 0$. It holds that $s_{N}\left({\bf p}|{\bf q}\right)=0$ if and only if ${\bf p}={\bf q}$.
@@
$p_{j}(y;z):=\left(G_{i}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(y)\right)_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}},\ j\geq 1.$
@@
$p_{0}(y;z):={\bf G}(y)\in\overline{\mathcal{P}_{N}}.$
By a martingale argument, we can show that
@@
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}(p_{j}(0;z)|{\bf q}),\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
@@
For $\epsilon\in(0,1)$, let $\overline{\mathcal{P}_{N}(\epsilon)}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq\epsilon,\sum_{i}p_{i}=1\right\}$. Then, $\overline{\mathcal{P}_{N}(\epsilon)}\subset\mathcal{P}_{N}\subset\overline{\mathcal{P}_{N}}$.
Recall Lemma 3.4. If
@@
then, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. In other words, ${\bf\widetilde{q}}$ is apart from the boundary $\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$
Assume that
@@
If |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$, then, there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
@@
We first give some further definitions for quantities of the relative entropy.
@@
Let
$V\left({\bf q};\epsilon\right):=\max\left\{s_{N}\left({\bf p}|{\bf q}\right)\middle|{\bf p}\in\overline{\mathcal{P}_{N}},\|{\bf p}-{\bf q}\|\geq\epsilon\right\},\ \epsilon>0.$
@@
(i) $V\left({\bf q};\epsilon\right)<0$.
(ii) $V\left({\bf q};\epsilon\right)$ is decreasing with respect to $\epsilon$.
(iii) $V\left({\bf q};\epsilon\right)$ is right-continuous with respect to $\epsilon$.
(iv) $V\left({\bf q};\epsilon\right)$ is continuous with respect to ${\bf q}$.
@@
(i) and (ii) are easy to see. (iii) Assume that for some $\epsilon_{0}\in(0,1/3)$, $V\left({\bf q};\epsilon\right)$ is not right-continuous. Then, $V\left({\bf q};\epsilon_{0}\right)>\lim_{\epsilon\to\epsilon_{0}+0}V\left({\bf q};\epsilon\right)$. Assume that $V\left({\bf q};\epsilon_{0}\right)=s_{N}\left({\bf p}|{\bf q}\right)$. If $\|{\bf p}-{\bf q}\|>\epsilon_{0}$, then, this cannot occur. Assume $\|{\bf p}-{\bf q}\|=\epsilon_{0}$. Then, there exists a sequence $\{{\bf p}^{(n)}\}_{n}$ in $\overline{\mathcal{P}_{N}}$ such that $\|{\bf p}^{(n)}-{\bf q}\|\geq\epsilon_{0}+N^{-n}$ for each $n$ and $\lim_{n\to\infty}\|{\bf p}^{(n)}-{\bf q}\|=\epsilon_{0}$. Hence,
@@
$V\left({\bf q};\epsilon+\|{\bf p}-{\bf q}\|\right)\leq V\left({\bf p};\epsilon\right)$
By Lemma 3.9, we can let
$V(\epsilon):=\max_{{\bf q}\in\overline{\mathcal{P}_{N}}}V\left({\bf q};\epsilon\right)<0.$
We now give a quantity defined by $Y,\{G_{i}\}_{i}$ and $\{H_{i}\}_{i}$.
@@
$W_{i}\left({\bf q},y\right):=\max\left\{\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|,\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|\right\},\ {\bf q}\in\overline{\mathcal{P}_{N}},y\in Y,i\in\mathcal{I}_{N}.$
@@
This assertion is easily shown. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left({\bf G}(y)|{\bf q}\right)<0$. However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$. Assume that |$\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|$ is close to 00, then, $\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|$ is strictly positive. Hence, $W_{i}\left({\bf q},y\right)>0$.
The following corresponds to [27, Lemma 8]. Informally speaking, if $W_{i}({\bf q},y)$ is positive uniformly to $y$, then, a sum of two successive terms of $s_{N}$ is negative.
@@
Let ${\bf q}\in\mathcal{P}_{N}$. Let $i_{1}\in\mathcal{I}_{N}$. Assume that for some $\epsilon_{1}>0$,
$\inf_{y\in Y}W_{i_{1}}\left({\bf q},y\right)\geq\epsilon_{1}.$ (3.8)
Assume that $X_{j+1}(z)=i_{1}$. Then,
$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)\leq V(\epsilon_{1})<0.$
@@
The proof is essentially the same as the proof of [27, Lemma 8] [^3] It contains a typo.. Since $X_{j+1}(z)=i_{1}$,
$p_{j}(0;z)={\bf G}\left(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right),$
@@
$p_{j+1}(0;z)={\bf G}\left(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right).$
@@
$\left|G_{k_{1}}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{1}}\right|\geq\epsilon_{1},\textup{ for some }k_{1}\in\mathcal{I}_{N},$
or,
$\left|G_{k_{2}}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{2}}\right|\geq\epsilon_{1},\textup{ for some }k_{2}\in\mathcal{I}_{N},$
holds. If the former holds, then, $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. If the latter holds, then, $s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. Now it suffices to recall that $s_{N}$ takes non-positive values. ∎
Let :=|$\ell:=|\{\mathfrak{p}_{i}\}_{i}|\geq 2$. We assume that $\{\mathfrak{p}_{i}\}_{i}=\{\mathfrak{p}_{j_{k}}:1\leq k\leq\ell\}$ where $10\leq j_{1}<\cdots<j_{\ell}\leq N-$ 1. Let $\delta_{0}:=\min\{\|\mathfrak{p}_{j_{k_{1}}}-\mathfrak{p}_{j_{k_{2}}}\|:k_{1}\neq k_{2}\}>0$. Let $B_{j}:=\left\{{\bf p}\in\overline{\mathcal{P}_{N}}:\|{\bf p}-\mathfrak{p}_{j}\|\leq\delta_{0}/4\right\}$. $B_{j}=B_{k}$ can occur even if $j\neq k$. Define a coloring map $\mathscr{A}:\overline{\mathcal{P}_{N}}\to\mathcal{I}_{N}$ by
$\mathscr{A}\left({\bf p}\right):=\begin{cases}j_{2}&{\bf p}\in B_{j_{1}}\\ \vdots&\\ j_{\ell}&{\bf p}\in B_{j_{\ell-1}}\\ j_{1}&{\bf p}\in B_{j_{\ell}}\\ 0&\textup{ otherwise}\end{cases}.$
Under these definitions, we see that
@@
$\widetilde{\epsilon_{1}}:=\inf\left\{W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\middle|{\bf q}\in\overline{\mathcal{P}_{N}},y\in Y\right\}>0.$
@@
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
If ${\bf G}(y)\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}={\bf G}\left(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})\right)$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$. Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$. Hence,
$\inf_{{\bf G}(y)\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
$\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
Assume that ${\bf G}(y)\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{N}\sum_{j}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{4N}.$
@@
$\inf_{y\in Y}W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\min\left\{\frac{\delta_{0}}{4N},\epsilon_{j_{k}}\middle|1\leq k\leq\ell\right\}=:\tilde{\epsilon_{1}}.$
@@
Let $I\subset\mathcal{I}_{N}$ be a non-empty set. Let $c_{I}:=\inf_{y\in Y}\sum_{i\in I}G_{i}(y)$. Then,
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
@@
If $c_{I}=0$, then, the assertion is obvious. Assume $c_{I}>0$. We denote the expectation with respect to $\nu_{0}$ by $E^{\nu_{0}}$. Let $\mathcal{F}_{m}:=\sigma(\xi_{1},\cdots,\xi_{m})$ for $m\geq 1$ and $\mathcal{F}_{0}$ be the trivial $\sigma$-algebra on $(\mathcal{I}_{N})^{\mathbb{N}}$. We denote the indicator function of $A\subset(\mathcal{I}_{N})^{\mathbb{N}}$ by ${\bf 1}_{A}$.
Let $M_{0}:=0$ and
$M_{n}-M_{n-1}:={\bf 1}_{\{\xi_{n}\in I\}}-E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right],n\geq 1.$
Then, $\{M_{n},\mathcal{F}_{n}\}_{n}$ is a martingale.
By (3.1), $E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right]\geq c_{I}$, $\nu_{0}$-a.s. Therefore, for every $\eta\in(0,1)$,
$\nu_{0}\left(\frac{|\{1\leq j\leq n:\xi_{j}\in I\}|}{n}\leq(1-\eta)c_{I}\right)\leq\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right).$
We remark that $|M_{n}-M_{n-1}|\leq 1$ for each $n\geq 1$. By Azuma’s inequality ([1], [31, E14.2]),

### v4

@@ 3.1. Proofs
$s_{N}\left({\bf p}|{\bf q}\right)\coloneqq\sum_{i\in\mathcal{I}_{N}}p_{i}\log\frac{q_{i}}{p_{i}},\ \ {\bf p}=(p_{i})_{i\in\mathcal{I}_{N}},{\bf q}=(q_{i})_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}}.$
If $p_{i}=0$, then we let $p_{i}\log\dfrac{q_{i}}{p_{i}}\coloneqq 0$. This is also valid for $q_{i}=0$. For each ${\bf q}\in\mathcal{P}_{N}$, the function $s_{N}(\cdot|{\bf q})$ is continuous on $\overline{\mathcal{P}_{N}}$. If ${\bf q}\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$, $s_{N}$ can take -$\infty$. By Jensen’s inequality, $s_{N}\left({\bf p}|{\bf q}\right)\leq 0$. It holds that $s_{N}\left({\bf p}|{\bf q}\right)=0$ if and only if ${\bf p}={\bf q}$.
@@
$p_{j}(y;z)\coloneqq\left(G_{i}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(y)\right)_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}},\ j\geq 1.$
@@
$p_{0}(y;z)\coloneqq{\bf G}(y)\in\overline{\mathcal{P}_{N}}.$
By a martingale argument, we can show the following assertion:
@@
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}(p_{j}(0;z)|{\bf q}),\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
@@
For $\epsilon\in(0,1)$, let
$\overline{\mathcal{P}_{N}(\epsilon)}\coloneqq\left\{(p_{i})_{i\in\mathcal{I}_{N}}\colon p_{i}\geq\epsilon,\sum_{i\in\mathcal{I}_{N}}p_{i}=1\right\}.$
Then, $\overline{\mathcal{P}_{N}(\epsilon)}\subset\mathcal{P}_{N}\subset\overline{\mathcal{P}_{N}}$.
Let $\widetilde{q}_{i}\coloneqq r_{i}^{s}>0$ and ${\bf\widetilde{q}}\coloneqq(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. Recall Lemma 3.4. If
@@
then, by (3.5), $\displaystyle r_{i}^{s}\geq\epsilon^{s}\geq\epsilon^{1+\frac{\log 2}{\log(1/\epsilon)}}=\dfrac{\epsilon}{2}$ and hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. In other words, ${\bf\widetilde{q}}$ has positive distance from the boundary $\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
Assume the following assertion:
@@
If $\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$, then there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\epsilon_{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
@@
We first give some further definitions for quantities of the relative entropy. We remark that the diameter of $\overline{\mathcal{P}}_{N}$ with respect to $\|\cdot\|_{1}$ is equal to $2$. By the triangle inequality, it holds that for ${\bf p}=(p_{i})_{i\in\mathcal{I}_{N}},{\bf q}=(q_{i})_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}}_{N}$,
$\left\|{\bf p}-{\bf q}\right\|_{1}=\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|\leq\sum_{i\in\mathcal{I}_{N}}(p_{i}+q_{i})=2.$
If ${\bf p}=(1,0,\dots,0)$ and ${\bf q}=(0,\dots,0,1)$, then $\left\|{\bf p}-{\bf q}\right\|_{1}=2$.
@@
For ${\bf q}\in\mathcal{P}_{N}$, let
$V\left({\bf q};\epsilon^{\prime}\right)\coloneqq\max\left\{s_{N}\left({\bf p}|{\bf q}\right)\middle|{\bf p}\in\overline{\mathcal{P}_{N}},\|{\bf p}-{\bf q}\|_{1}\geq\epsilon^{\prime}\right\},\ \epsilon^{\prime}\in(0,1].$
@@
(i) $V\left({\bf q};\epsilon^{\prime}\right)<0$.
(ii) $V\left({\bf q};\epsilon^{\prime}\right)$ is decreasing with respect to $\epsilon^{\prime}$.
(iii) $V\left({\bf q};\epsilon^{\prime}\right)$ is right-continuous with respect to $\epsilon^{\prime}$.
(iv) $V\left({\bf q};\epsilon^{\prime}\right)$ is continuous with respect to ${\bf q}$.
@@
Assertions (i) and (ii) are easy to see. (iii) Assume that for some $\epsilon_{0}\in(0,1/3)$, $V\left({\bf q};\epsilon^{\prime}\right)$ is not right-continuous. Then, $\displaystyle V\left({\bf q};\epsilon_{0}\right)>\lim_{\epsilon^{\prime}\to\epsilon_{0}+0}V\left({\bf q};\epsilon^{\prime}\right)$. Assume that $V\left({\bf q};\epsilon_{0}\right)=s_{N}\left({\bf p}|{\bf q}\right)$. If $\left\|{\bf p}-{\bf q}\right\|_{1}>\epsilon_{0}$, then this cannot occur. Assume $\left\|{\bf p}-{\bf q}\right\|_{1}=\epsilon_{0}$. Then, there exists a sequence $\{{\bf p}^{(n)}\}_{n}$ in $\overline{\mathcal{P}_{N}}$ such that $\left\|{\bf p}^{(n)}-{\bf q}\right\|_{1}\geq\epsilon_{0}+N^{-n}$ for each $n$ and $\displaystyle\lim_{n\to\infty}\left\|{\bf p}^{(n)}-{\bf q}\right\|_{1}=\epsilon_{0}$. Hence,
@@
$V\left({\bf q};\epsilon^{\prime}+\left\|{\bf p}-{\bf q}\right\|_{1}\right)\leq V\left({\bf p};\epsilon^{\prime}\right)\qed$
Setting
$V_{\epsilon,1}(\epsilon^{\prime})\coloneqq\max_{{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}}V\left({\bf q};\epsilon^{\prime}\right),\ \ \epsilon^{\prime}\in(0,1],$
we obtain by Lemma 3.9 and the compactness of $\overline{\mathcal{P}_{N}(\epsilon/2)}$ that
$V_{\epsilon,1}(\epsilon^{\prime})<0.$
@@
$W_{i}\left({\bf q},y\right)\coloneqq\max\left\{\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|,\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|\right\},\ {\bf q}\in\overline{\mathcal{P}_{N}},y\in Y,i\in\mathcal{I}_{N}.$
@@
This assertion is easily shown and we omit the proof.
If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then $s_{N}\left({\bf G}(y)|{\bf q}\right)<0$. However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$. Assume that $\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. If $\displaystyle\max_{k\in\mathcal{I}_{N}}\left|G_{k}(y)-q_{k}\right|$ is close to $0$, then $\displaystyle\max_{k\in\mathcal{I}_{N}}\left|G_{k}(H_{i}(y))-q_{k}\right|$ is strictly positive. Hence, $W_{i}\left({\bf q},y\right)>0$.
Since the diameter of $\overline{\mathcal{P}}_{N}$ is equal to $2$, we obtain $0\leq W_{i}\left({\bf q},y\right)\leq 2$.
The following corresponds to [27, Lemma 8]. Informally speaking, if $W_{i}({\bf q},y)$ is positive uniformly in $y$, then a sum of two successive terms of $s_{N}$ is negative.
@@
Let ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Let $i_{1}\in\mathcal{I}_{N}$. Assume that for some $\epsilon_{1}\in(0,1]$,
(3.9) $\inf_{y\in Y}W_{i_{1}}\left({\bf q},y\right)\geq\epsilon_{1}.$
Assume that $z\in\left(\mathcal{I}_{N}\right)^{\mathbb{N}}$ and $\xi_{j+1}(z)=i_{1}$. Then,
$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)\leq V_{\epsilon,1}(\epsilon_{1})<0.$
@@
The proof is essentially the same as the proof of [27, Lemma 8] [^3] The proof contains a typo. In Case 2 in the proof, $G_{i+l}(h_{i+l}(y;x))$ should be replaced with $G_{i_{l}}(h_{i+l}(y;x))$.. Since $\xi_{j+1}(z)=i_{1}$,
$p_{j}(0;z)={\bf G}\left(H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0)\right),$
@@
$p_{j+1}(0;z)={\bf G}\left(H_{i_{1}}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0)\right).$
@@
$\left|G_{k_{1}}(H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0))-q_{k_{1}}\right|\geq\epsilon_{1},\textup{ for some }k_{1}\in\mathcal{I}_{N}$
or
$\left|G_{k_{2}}(H_{i_{1}}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0))-q_{k_{2}}\right|\geq\epsilon_{1},\textup{ for some }k_{2}\in\mathcal{I}_{N},$
holds. We remark that at least one of them holds. If the former holds, then $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. If the latter holds, then $s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. Now it suffices to recall that $s_{N}$ takes non-positive values. ∎
Let $\ell\coloneqq\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. We assume that $\{\mathfrak{p}_{i}\}_{i}=\{\mathfrak{p}_{j_{k}}\colon 1\leq k\leq\ell\}$ where $0\leq j_{1}<\cdots<j_{\ell}\leq N-1$. Let
$\displaystyle\delta_{0}$ $\displaystyle\coloneqq\min\left\{\left\|\mathfrak{p}_{j_{k_{1}}}-\mathfrak{p}_{j_{k_{2}}}\right\|_{1}\colon k_{1}\neq k_{2}\right\}>0,\ \ \textup{ and }$
$\displaystyle B_{j}$ $\displaystyle\coloneqq\left\{{\bf p}\in\overline{\mathcal{P}_{N}}\colon\left\|{\bf p}-\mathfrak{p}_{j}\right\|_{1}\leq\delta_{0}/4\right\}.$
The equality $B_{j}=B_{k}$ can occur even if $j\neq k$. Define a coloring map $\mathscr{A}\colon\overline{\mathcal{P}_{N}}\to\mathcal{I}_{N}$ by
$\mathscr{A}\left({\bf p}\right)\coloneqq\begin{cases}j_{2}&{\bf p}\in B_{j_{1}}\\ \vdots&\\ j_{\ell}&{\bf p}\in B_{j_{\ell-1}}\\ j_{1}&{\bf p}\in B_{j_{\ell}}\\ 0&\textup{ otherwise}\end{cases}.$
Under these definitions, we see the following assertion:
@@
$\widetilde{\epsilon_{1}}\coloneqq\inf\left\{W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\middle|{\bf q}\in\overline{\mathcal{P}_{N}},y\in Y\right\}>0.$
@@
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
Let
$\widetilde{B}_{j}\coloneqq\left\{{\bf p}\in\overline{\mathcal{P}_{N}}\colon\left\|{\bf p}-\mathfrak{p}_{j}\right\|_{1}\leq\delta_{0}/8\right\},\ j\in\mathcal{I}_{N}.$
If ${\bf G}(y)\notin\widetilde{B}_{\mathscr{A}({\bf q})}$, then $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}={\bf G}\left(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})\right)$ and hence, there exists an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$ but not $y$. Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\displaystyle\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$. Hence,
$\inf_{{\bf G}(y)\notin\widetilde{B}_{\mathscr{A}({\bf q})}}\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
$\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
Assume that ${\bf G}(y)\in\widetilde{B}_{\mathscr{A}({\bf q})}$. First, we consider the case where $\displaystyle{\bf q}\in\bigcup_{j\in\mathcal{I}_{N}}B_{j}$. Then, there exists a unique $i\in\{1,\dots,\ell\}$ such that ${\bf q}\in B_{j_{i}}$ and $\mathscr{A}({\bf q})\neq j_{i}$. By the triangle inequality,
$\left\|G(y)-{\bf q}\right\|_{1}\geq\left\|\mathfrak{p}_{\mathscr{A}({\bf q})}-\mathfrak{p}_{j_{i}}\right\|_{1}-\left\|G(y)-\mathfrak{p}_{\mathscr{A}({\bf q})}\right\|_{1}-\left\|{\bf q}-\mathfrak{p}_{j_{i}}\right\|_{1}\geq\frac{\delta_{0}}{2}.$
Second, we consider the case where $\displaystyle{\bf q}\notin\bigcup_{j\in\mathcal{I}_{N}}B_{j}$. Then, ${\bf q}\notin B_{\mathscr{A}({\bf q})}$. By the triangle inequality,
$\left\|G(y)-{\bf q}\right\|_{1}\geq\left\|\mathfrak{p}_{\mathscr{A}({\bf q})}-{\bf q}\right\|_{1}-\left\|\mathfrak{p}_{\mathscr{A}({\bf q})}-G(y)\right\|_{1}\geq\frac{\delta_{0}}{8}.$
Thus we obtain that
$W_{\mathscr{A}\left({\bf q}\right)}\left({\bf q},y\right)\geq\frac{1}{N}\sum_{j\in\mathcal{I}_{N}}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{8N}.$
@@
$\inf_{y\in Y}W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\min\left\{\frac{\delta_{0}}{8N},\epsilon_{j_{k}}\middle|1\leq k\leq\ell\right\}=:\widetilde{\epsilon_{1}}$$\qed$
@@
Let $I\subset\mathcal{I}_{N}$ be a non-empty set. Let $\displaystyle c_{I}\coloneqq\inf_{y\in Y}\sum_{i\in I}G_{i}(y)$. Then,
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n\colon\xi_{j}(z)\in I\}|}{n}\geq c_{I},\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
@@
If $c_{I}=0$, then the assertion is obvious. Assume $c_{I}>0$. We regard $\{\xi_{j}\}_{j\geq 1}$ as random variables with respect to the probability measure $\nu_{0}$. We denote the expectation with respect to $\nu_{0}$ by $E^{\nu_{0}}$. Let $\mathcal{F}_{m}\coloneqq\sigma(\xi_{1},\ldots,\xi_{m})$ for $m\geq 1$ and $\mathcal{F}_{0}$ be the trivial $\sigma$-algebra on $(\mathcal{I}_{N})^{\mathbb{N}}$. We denote the indicator function of $A\subset(\mathcal{I}_{N})^{\mathbb{N}}$ by ${\bf 1}_{A}$.
Let $M_{0}\coloneqq 0$ and
$M_{j}-M_{j-1}\coloneqq{\bf 1}_{\{\xi_{j}\in I\}}-E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{j}\in I\}}|\mathcal{F}_{j-1}\right],\ j\geq 1.$
Then, $\left\{M_{n},\mathcal{F}_{n}\right\}_{n}$ is a martingale.
By (3.1), for every $j\geq 1$, $E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{j}\in I\}}|\mathcal{F}_{j-1}\right]\geq c_{I}$, $\nu_{0}$-a.s. Then, we see that
$\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{I}(\xi_{i})=\frac{M_{n}}{n}+\frac{1}{n}\sum_{i=1}^{n}E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{j}\in I\}}|\mathcal{F}_{j-1}\right]\geq\frac{M_{n}}{n}+c_{I},\textup{ $\displaystyle\nu_{0}$-a.s.}$
Therefore, it holds that for every $\eta\in(0,1)$,
$\displaystyle\nu_{0}\left(\frac{|\{1\leq j\leq n\colon\xi_{j}\in I\}|}{n}\leq(1-\eta)c_{I}\right)$ $\displaystyle\leq\nu_{0}\left(\frac{M_{n}}{n}+c_{I}\leq(1-\eta)c_{I}\right)$
$\displaystyle=\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right).$
We remark that $|M_{n}-M_{n-1}|\leq 1$ for each $n\geq 1$. By Azuma’s inequality ([2], [33, E14.2]),